4. x^2 + 3x = 28


5.2x^2 + 9x - 5 = 0

Answer by stanbon(75887)

4. x^2 + 3x = 28
x^2 + 3x - 28 = 0
(x+7)(x-4) = 0
x = -7 or x = 4
============================

5.2x^2 + 9x - 5 = 0
2x^2 + 10x - x -5 = 0
2x(x+5) -(x+5) = 0
(x+5(2x-1) = 0
x = -5 or x = 1/2
========================
Cheers,
Stan H.